You want to buy a new sports coupe for $75,500, and the finance office at the dealership has quoted you a 5.7 percent APR loan for 60 months to buy the car.

What will your monthly payments be? (Do not round intermediate calculations and round your final answer to 2 decimal places. (e.g., 32.16))


  Monthly payment $   

What is the effective annual rate on this loan? (Do not round intermediate calculations and round your final answer to 2 decimal places. (e.g., 32.16))


  Effective annual rate %  
Answer #1

Hi,


Please find the answer as follows:



Part A:


Nper = 60 (indicates the period over which the payments are made)

PV = 75500 (indicates the value of loan)

Rate = 5.7%/12 (indicates monthly interest rate)

FV = 0 (indicates future value of loan, not relevant in this case)

PMT = ? (indicates monthly payment)


Amount of Monthly Payment = PMT(Rate,Nper,PV,FV) = PMT(5.7%/12,60,-75500,0) = 1449.12


Part B:


Effective Annual Rate of Interest = (1+R/12)^12 - 1 = (1+5.7%/12)^12 - 1 = 5.85%


Thanks.
Answer #2

Hi,


Please find the answer as follows:



Part A:


Nper = 60 (indicates the period over which the payments are made)

PV = 75500 (indicates the value of loan)

Rate = 5.7%/12 (indicates monthly interest rate)

FV = 0 (indicates future value of loan, not relevant in this case)

PMT = ? (indicates monthly payment)


Amount of Monthly Payment = PMT(Rate,Nper,PV,FV) = PMT(5.7%/12,60,-75500,0) = 1449.12


Part B:


Effective Annual Rate of Interest = (1+R/12)^12 - 1 = (1+5.7%/12)^12 - 1 = 5.85%


Thanks.